OpenStudy (anonymous):

A turtle(neithermutated nor ninja) starts crawling from rest. This ordinary turtle accelerates at 6E-4 m/s^2. After 30 seconds at this strenuous rate, find it's velocity and the distance covered

OpenStudy (anonymous):

\[s_x=\frac{1}{2}at^2+v_{0x}t+s_0\] \[s_x=\frac{1}{2}(.0006)(30^2)+0+0=.27 m\]

OpenStudy (anonymous):

\[v_x=at+v_0\] \[v_x=(.0006)(30)=.018m/s^2\]

OpenStudy (anonymous):

In order to catch a speeder, a police officer must achieve a speed of 40m/s within a distance of 1km. Presuming the officer tarts from rest, how much acceleration is required?

OpenStudy (anonymous):

\[1km=1000m\] \[s_0=\frac{1}{2}at^2 +v_0 t+s_0\] \[s_0=\frac{1}{2}at\]

OpenStudy (anonymous):

\[1000=\frac{1}{2}at^2\]

OpenStudy (anonymous):

\[v_x=at+0\] \[40=at\] \[t=\frac{40}{a}\] \[1000=\frac{1}{2}a(\frac{40}{a})^2 \]

OpenStudy (anonymous):

\[1000=.5(\frac{40^2}{a})\] \[a=\frac{.5(40^2)}{1000}\] \[a=.8m/s^2\]